what is the final pressure of an ideal gas at 50 psig and 90 F, heated in a closed container to 150 F?expres your answer in psia
1
Expert's answer
2020-11-09T14:17:52-0500

Absolute pressure = Gauge pressure - atmospheric pressure.

"T_1=90F =\frac{(90-32)*5}{9}+273=305 K"

"P_1=50 psig =(50+14.5) = 64.5 psi absolute"

"T_2=150F =\frac{(150-32)*5}{9}+273=338.7 K"

"\frac{P_1}{T_1}=\frac{P_2}{T_2}
\implies P_2=\frac{P_1*T_2}{T_1}"

"P_2=\frac{338.7*64.5}{305.4}=71.53 abs=71.53 psia"